On continuous solutions of some functional equations Janusz Brzdȩk, Rzeszów, Poland

Keywords: functional equation, continuous function, associative binary operation AMS Subject classification: 39B22 Abstract: There is given a method of finding continuous solutions f of some functional equations of form (1) in several classes of functions mapping a real interval into the set of reals.